# Quantum Critical Behavior in Ce-Au-Al Quasicrystal Approximants

### v1

@@ I Introduction
Understanding the physical properties of aperiodic quasicrystals (QCs) is challenging, since QCs cannot be easily treated with conventional methods Levine and Steinhardt (1984, 1986); Goldman and Widom (1991); Tsai (1999). The structurally and compositionally closely related approximant crystals (ACs), however, are built of the same atomic clusters as QCs but possess periodicity in 3D. Experimentally observed are 1/1 and higher order 2/1 ACs. Especially in the latter, the spacial arrangement of cluster building units connects closely to QCs Takakura et al. (2007). An important question is whether QCs possess specific physical properties that are exclusive to their long-range quasiperiod order Wessel et al. (2003); Goldman et al. (2013); Goldman (2014); Thiem and Chalker (2015); Takeuchi et al. (2023). In this respect, quantum critical behavior has been reported in the icosahedral QC Yb-Au-Al (Yb15Au51Al34) Deguchi et al. (2012); Watanuki et al. (2012), which shows a divergence of several physical properties like magnetic susceptibility and specific heat, when the temperature is decreased. In contrast, the critical fluctuations are suppressed at low temperatures in the 1/1 AC (Yb14Au51Al35) Deguchi et al. (2012). Thus, aperiodicity brings the Yb-Au-Al system closer to the quantum critical point. The question, however, remains whether quasicrystals are fundamentally different from their approximants in this regard.
The Yb-Au-Al intermetallic system contains a QC and a 1/1 AC phase. Here, the focus is on the Ce-Au-Al system which hosts 2/1 and 1/1 AC phases, whereas the QC phase is absent Muro et al. (2021). Muro et al. Muro et al. (2021) have reported on 1/1 AC and 2/1 AC specimens with various Au concentrations in the Ce-Au-Al system. From their study of zero-field specific heat, magnetic susceptibility at 1 T, and resistivity, they concluded that these systems show signatures of a Kondo resonance as well as spin-glass behavior below 1 K. Here, we study the effect of magnetic field on the ac-susceptibility and low-temperature specific heat of single crystal 1/1 AC and 2/1 AC samples. The ac-susceptibility measurements show a diverging behavior for low temperatures, indicating quantum critical fluctuations even in these approximants, while the low-temperature specific heat is dominated by a magnetic Schottky anomaly due to a splitting of the Kramers doublet ground state of Ce3+.
@@ II Experimental Details
The samples by Muro et al. Muro et al. (2021) were synthesized using arc melting with stoichiometric compositions Ce14AuxAl86-x and Ce16AuxAl84-x, followed by annealing at 800 °C for 50 hours. Although homogeneous and polycrystalline samples were produced, an inseparable secondary phase, the hexagonal Ce14Au51 compound Trovarelli et al. (1994), coexisted in most of the samples. This secondary phase potentially contributed to the magnetic observations of the specimens. Therefore, to avoid the formation of the hexagonal Ce14Au51 compound, we applied a different synthesis approach aimed at isolating phase-pure, well-faceted single crystals with selected compositions of Ce14Au71Al15 for 1/1 AC and Ce12Au65Al23 for 2/1 AC.
Figure 1: Motifs of the cubic 1/1 and 2/1 AC structures in the Ce-Au-Al system. (a). Polyhedron shells of the Tsai cluster which is the fundamental unit of both structures. RTH: rhombic triacontahedron, ICD: icosidodecahedron, ICS: icosahedron, DO: dodecahedron, TE: tetrahedron. The ICS shell is exclusively built from Ce atoms, indicated in yellow. Grey circles signify pure Au, Al, or mixed Au/Al positions found throughout the various other shells. (b). Spatial distribution of ICS shells within the unit cells of the 1/1 and 2/1 ACs. In the 2/1 AC, which more closely relates to QC, an extra Ce atomic position is observed outside the ICS within the acute rhombohedron (AR) unit, marked by black lines, which fills the space between the RTH shells Gómez and Lidin (2001). (c) and (d). Symmetry-independent Ce positions in the 2/1 AC, situated within both the ICS and the double Friauf (DF) polyhedral inside the AR. The various colored spheres in (d) represent distinct Ce atoms, with Ce1-4 in the ICS, and Ce5 in the DF. The ICS’s twelve vertices suggest twelve pseudo-5-fold rotational axes (5f) throughout the structure. Each vertex’s Ce atom is central within a 16-atom polyhedron, bridging the inner DO shell to the outer RTH shell along the pseudo 5f axis, as shown in (c).
@@
Structural characterization was performed using powder and single-crystal X-ray diffraction. For compositional analysis, energy-dispersive X-ray spectroscopy on cross-section polished samples was utilized using a Zeiss LEO 1550 scanning electron microscope (for details see Supplementary Fig. S1). The ac-susceptibility was measured using a PPMS (Quantum Design Dynacool 9 T ACMS II). Specific heat measurements were performed in a Bluefors dilution refrigerator using a membrane-based nanocalorimeter Tagliati et al. (2012); Willa et al. (2017) with a sample size of 20 nmol for 1/1 AC and 15 nmol for 2/1 AC. Measurements were performed in the ac-steady state mode with continuously adjusted frequency for good accuracy, differential thermometry running at the 5th harmonic of the temperature oscillation, and local temperature control on the calorimeter membrane with the dilution refrigerator sitting at base temperature.
@@ III Results and Discussion
The structures of Ce-Au-Al ACs are built of so-called Tsai clusters which are the fundamental structural unit of icosahedral Tsai-type QCs Tsai et al. (2000). A Tsai cluster features a multi-shell arrangement of concentric polyhedra with the outermost shell being a rhombic triacontahedron (RTH), followed by an icosidodecahedron (ICD), an icosahedron (ICS), a dodecahedron (DO), and a tetrahedron (TE) at the center (Fig. 1a). In terms of crystallography, both 1/1 and 2/1 ACs are periodic crystals in 3D space, each with a distinct cubic space group, which leads to different Tsai-cluster arrangements in their unit cells (Fig. 1b) Gómez and Lidin (2001); Takakura et al. (2007). The 1/1 AC typically forms a body-centered cubic (bcc) structure with the space group Im-3 (However, we noticed superstructure reflections for the Ce-Au-Al 1/1 AC which suggest a lower symmetry, see Supplementary Fig. S1). In contrast, the 2/1 AC adopts a primitive cubic structure with the space group Pa-3. The 16-atom polyhedron, constructed from the vertices of the ICS along the pseudo-5f axis, is considered the coordination environment for the Ce atoms (Fig. 1c,d).
@@
At high temperatures, above $\sim 10$ K the behavior is dominated by the phonons. At the lowest temperatures, below 0.3 K, the upturn in $C/T$ is due to a nuclear specific heat arising from 197Au with nuclear spin $I=3/2$ and 27Al with $I=5/2$ Stone (2016), apparent at high magnetic fields. The dominant contribution of the low-temperature specific heat resembles a magnetic Schottky anomaly, shifting up in temperature with increasing magnetic field.
@@
Figure 3: Specific heat and entropy of 1/1 AC and 2/1 AC. (a) and (c). The specific heat after removing the phonon contribution (see Supplementary Fig. S2 and S3) is shown as $C/T$ for 1/1 and 2/1 AC, respectively. (b) and (d). Corresponding entropy of the systems. The dotted gray line indicates the value $R$ln2. (e). Sketch of splitting of Ce3+ ($J$=5/2) degenerate levels, due to the non-cubic crystalline field, into three Kramers doublets. Crystal field disorder or magnetic interactions in the system could be the reason for lifting the Kramers spin degeneracy (LKSD). The magnetic field further splits the ground state doublet $|0\rangle$ as seen in the panel e.
As we apply the magnetic field the peak in $C/T$ shifts to higher temperature. This magnetic field behavior of the temperature dependence shows a maximum in $C(T)$ that takes approximately the same value for all applied magnetic fields (see Supplementary Fig. S3). This indicates that the behavior is originating from a Schottky anomaly. The corresponding entropies, obtained as the integral of $C/T$, are shown in Fig. 3b and d. As seen from the entropy analysis, the entropy saturates to a value close to $\sim R\ln 2$ for both ACs at around 10 K. This indicates that the Schottky anomaly arises from a two-level system.
The non-cubic crystal field splits the Ce3+ ($J=5/2$) ground state into three Kramers doublets Kramers (1930); Wigner (1993); Desgranges (2014); Griffith et al. (2019). A sketch of the energy level splitting is shown in Fig. 3e, where the ground state Kramers doublet is denoted as $|0\rangle$. The external magnetic field further splits the ground state doublet due to the Zeeman effect. This effect of the magnetic field on the Kramers doublet is seen in the specific heat in Fig. 3a and c. This ground state Kramers doublet, thus, determines the low-temperature behavior in specific heat.
Figure 4: Temperature and magnetic field dependence of the crossover behavior. (a) and (b). The ac-susceptibility ($\chi$) for 1/1 and 2/1 AC for various magnetic fields in the log-lin scale. The inset shows the inverse susceptibility for several magnetic fields at low temperatures. The dotted line corresponds to $1/\chi\propto T$. (c) and (d). Magnetic field dependence of ac-susceptibility ($\chi$) for several temperatures in log-log scale. The applied magnetic field cuts off this diverging behavior and $\chi$ starts to decrease above the cut-off field ($B_{\alpha}$), shown as arrows. (e) and (f). Contour plot of $B$ vs. $T$ phase diagram for the ac-susceptibility $\chi$ showing the temperature-field dependence of the crossover behavior extracted from Panels a and b. The dashed line indicates the crossover region.
@@
The presence of a significant zero-field specific heat indicates that the Kramers doublet ground state is not degenerate in the absence of magnetic field. This lifting of Kramers degeneracy could result from crystal field disorder Krempaskỳ et al. (2024) or magnetic interactions within the system Winkler et al. (2003); Šmejkal et al. (2022a, b); Lee et al. (2024); Krempaskỳ et al. (2024).
@@
${W=\frac{\mu_{0}N_{\mathrm{A}}g_{\mathrm{eff}}^{2}\mu_{\mathrm{B}}^{2}}{3k_{\mathrm{B}}}}$
@@
is the Curie constant. The Curie temperature extracted from a fit of $1/\chi$ vs $T$, shown in Supplementary Fig. S4, is $-10\pm 0.5\,\mathrm{K}$ and $-6\pm 1\,\mathrm{K}$ for 1/1 and 2/1 ACs, respectively, indicating weakly antiferromagnetic coupling. The effective magnetic moment $\mu_{\mathrm{eff}}=g_{\mathrm{eff}}\mu_{\mathrm{B}}$ of Ce obtained from this high-temperature region is $\mu_{\mathrm{eff}}=2.53\pm 0.02\,\mu_{\mathrm{B}}$ and $\mu_{\mathrm{eff}}=2.50\pm 0.02\,\mu_{\mathrm{B}}$ for 1/1 and 2/1 ACs, respectively (see Supplementary Fig. S4). This effective magnetic moment is close to the moment for free Ce3+ ion ($2.54\,\mu_{\mathrm{B}}$).
@@
Going to lower temperatures, the zero-field susceptibility no longer follows the high-temperature Curie-Weiss law. The insets of Fig. 4a and b show the low-temperature inverse susceptibility. A Curie-Weiss fit for temperatures between 2-10 K yields $\mu_{\mathrm{eff}}=1.98\pm 0.02\,\mu_{\mathrm{B}}$ and $\mu_{\mathrm{eff}}=2.10\pm 0.02\,\mu_{\mathrm{B}}$ for 1/1 and 2/1 ACs, respectively (see Supplementary Fig. S4), with Curie temperatures well below 1 K. The low-temperature zero-field $\chi$, thus, shows essentially a power-law behavior ${\chi\propto 1/T}$. The fairly large $\mu_{\mathrm{eff}}$ implies that the ground state Kramers doublet has a significant contribution of $m_{J}=5/2$.
@@
The diverging behavior in $\chi(T)$ is suppressed when an external magnetic field is applied, as seen in Fig. 4a and b. Figure 4c and d show the magnetic field dependence of 1/1 and 2/1 AC, respectively, for several temperatures. For a given temperature there exists a cut-off field $B_{\alpha}$ above which this happens. Well below $B_{\alpha}$ the susceptibility becomes independent of the magnetic field and follows the zero-field $\chi(T)$. Above $B_{\alpha}$, the decrease of $\chi$ with field follows a power-law behavior. At the highest fields and lowest temperatures, the behavior becomes independent of temperature and the magnetic field controls the cut-off behavior.
The diverging zero-field susceptibility $\chi\propto 1/T$ could naively be attributed to a simple, non-interacting paramagnet with $\theta=0$. However, such a non-interacting system would not display any magnetic field dependence at low susceptibilities. The magnetic field cut-off behavior rather indicates that the observed behavior should be understood in terms of quantum criticality.
@@
For a quantum critical system, the susceptibility has a form of ${1/\chi\propto T^{\beta}}$, where $\beta$ is a critical exponent. This critical exponent characterizes the spectrum of critical fluctuations in a quantum critical system Löhneysen et al. (2007). For Ce-Au-Al ACs at low temperatures (2-10 K), this critical exponent seems to be approximately equal to one ($\beta=0.94\pm 0.01$ for 1/1 AC and $0.97\pm 0.01$ for 2/1 AC).
From transport measurements Muro et al. (2021), it is seen that the ACs of the Ce-Au-Al show a Kondo resonance peak at $T_{\mathrm{K}}\approx$1 K. This suggests that the local $f$-electrons are hybridized with the conduction electrons below $T_{\mathrm{K}}$. Even though $T_{\mathrm{K}}$ is low for these AC systems, the magnetic susceptibility shows the effects of strong interactions above this temperature. This might come from the Ruderman–Kittel–Kasuya–Yosida (RKKY) interactions (with energy scale $T_{\mathrm{RKKY}}$) driving the critical behavior observed in $\chi$. For a system where $T_{\mathrm{K}}<T_{\mathrm{RKKY}}$, the expected ground state could be either a long-range magnetic order or a spin-liquid state Coleman and Nevidomskyy (2010); Khansili et al. (2022).
Figure 4e and f illustrate the crossover behavior as a contour plot for both AC systems. The dashed line indicates the approximate location of the crossover region below which the system does not show critical behavior. From the temperature and field dependence of low-temperature $\chi$, it can be seen that $\chi$ shows a temperature-field competition that sets the cut-off scale in this region. This is one of the defining features of quantum criticality in a broad range of systems Millis et al. (2002); Löhneysen et al. (2007); Gegenwart et al. (2008); Matsumoto et al. (2011); Wu et al. (2018); Khansili et al. (2023) including the quantum critical quasicrystal Yb-Au-Al Deguchi et al. (2012); Watanuki et al. (2012); Shaginyan et al. (2013).
From the present data, it is not clear whether there exists an additional intrinsic scale that cuts off the critical behavior even in zero field, either by becoming magnetically ordered or by remaining as a quantum paramagnet sensitive to magnetic field. Such a cut-off would certainly be below 2 K as seen from Fig. 4. If magnetic ordering sets in, the transition temperature is below about 0.2 K, since no clear signatures are seen in specific heat. Possible enhancement of the conduction electron effective mass above 5 K as extracted from the Sommerfeld coefficient evaluated in Supplementary Fig. S2 is rather modest, with $\gamma\approx 4$ mJ/mol-at.K2. This indicates that the system does not display any significant heavy fermion characteristics, consistent with having $T_{\mathrm{K}}<T_{\mathrm{RKKY}}$ Coleman and Nevidomskyy (2010); Khansili et al. (2022).

### v2

# Precursor to Quantum Criticality in Ce-Au-Al Quasicrystal Approximants

@@ I Introduction
Understanding the physical properties of aperiodic quasicrystals (QCs) is challenging, since QCs cannot be easily treated with conventional methods [1, 2, 3, 4]. The structurally and compositionally closely related approximant crystals (ACs), however, are built of the same atomic clusters as QCs but possess periodicity in 3D. Experimentally observed are 1/1 and higher order 2/1 ACs. Especially in the latter, the spacial arrangement of cluster building units connects closely to QCs [5]. An important question is whether QCs possess specific physical properties that are exclusive to their long-range quasiperiod order [6, 7, 8, 9, 10]. In this respect, quantum critical behavior has been reported in the icosahedral QC Yb-Au-Al (Yb15Au51Al34) [11, 12], which shows a divergence of several physical properties like magnetic susceptibility and specific heat, when the temperature is decreased. In contrast, the critical fluctuations are suppressed at low temperatures in the 1/1 AC (Yb14Au51Al35) [11]. Thus, aperiodicity brings the Yb-Au-Al system closer to the quantum critical point. The question, however, remains whether quasicrystals are fundamentally different from their approximants in this regard.
The Yb-Au-Al intermetallic system contains a QC and a 1/1 AC phase. Here, the focus is on the Ce-Au-Al system which hosts 2/1 and 1/1 AC phases, whereas the QC phase is absent [13]. Muro et al. [13] have reported on 1/1 AC and 2/1 AC specimens with various Au concentrations in the Ce-Au-Al system. From their study of zero-field specific heat, magnetic susceptibility at 1 T, and resistivity, they concluded that these systems show signatures of a Kondo resonance as well as spin-glass behavior below 1 K. Here, we study the effect of magnetic field on the ac-susceptibility and low-temperature specific heat of single crystal 1/1 AC and 2/1 AC samples. The ac-susceptibility measurements show a diverging behavior for low temperatures, and the magnetic field behavior closely resembles that of other quantum critical systems. The low-temperature specific heat is dominated by a magnetic Schottky anomaly due to a splitting of the Kramers doublet ground state of Ce3+.
@@ II Experimental Details
The samples by Muro et al. [13] were synthesized using arc melting with stoichiometric compositions Ce14AuxAl86-x and Ce16AuxAl84-x, followed by annealing at 800 °C for 50 hours. Although homogeneous and polycrystalline samples were produced, an inseparable secondary phase, the hexagonal Ce14Au51 compound [14], coexisted in most of the samples. This secondary phase potentially contributed to the magnetic observations of the specimens. Therefore, to avoid the formation of the hexagonal Ce14Au51 compound, we applied a different synthesis approach aimed at isolating phase-pure, well-faceted single crystals with selected compositions of Ce14Au71Al15 for 1/1 AC and Ce12Au65Al23 for 2/1 AC.
Figure 1: Motifs of the cubic 1/1 and 2/1 AC structures in the Ce-Au-Al system. (a). Polyhedron shells of the Tsai cluster which is the fundamental unit of both structures. RTH: rhombic triacontahedron, ICD: icosidodecahedron, ICS: icosahedron, DO: dodecahedron, TE: tetrahedron. The ICS shell is exclusively built from Ce atoms, indicated in yellow. Grey circles signify pure Au, Al, or mixed Au/Al positions found throughout the various other shells. (b). Spatial distribution of ICS shells within the unit cells of the 1/1 and 2/1 ACs. In the 2/1 AC, which more closely relates to QC, an extra Ce atomic position is observed outside the ICS within the acute rhombohedron (AR) unit, marked by black lines, which fills the space between the RTH shells [15]. (c) and (d). Symmetry-independent Ce positions in the 2/1 AC, situated within both the ICS and the double Friauf (DF) polyhedral inside the AR. The various colored spheres in (d) represent distinct Ce atoms, with Ce1-4 in the ICS, and Ce5 in the DF. The ICS’s twelve vertices suggest twelve pseudo-5-fold rotational axes (5f) throughout the structure. Each vertex’s Ce atom is central within a 16-atom polyhedron, bridging the inner DO shell to the outer RTH shell along the pseudo 5f axis, as shown in (c).
@@
Structural characterization was performed using powder and single-crystal X-ray diffraction. For compositional analysis, energy-dispersive X-ray spectroscopy on cross-section polished samples was utilized using a Zeiss LEO 1550 scanning electron microscope (for details see Supplementary Fig. S1 and Fig. S2). The ac-susceptibility was measured using a PPMS (Quantum Design Dynacool 9 T ACMS II). Specific heat measurements were performed in a Bluefors dilution refrigerator using a membrane-based nanocalorimeter [18, 19] with a sample size of 20 nmol for 1/1 AC and 15 nmol for 2/1 AC. Measurements were performed in the ac-steady state mode with continuously adjusted frequency for good accuracy, differential thermometry running at the 5th harmonic of the temperature oscillation, and local temperature control on the calorimeter membrane with the dilution refrigerator sitting at base temperature.
@@ III Results and Discussion
The structures of Ce-Au-Al ACs are built of so-called Tsai clusters which are the fundamental structural unit of icosahedral Tsai-type QCs [20]. A Tsai cluster features a multi-shell arrangement of concentric polyhedra with the outermost shell being a rhombic triacontahedron (RTH), followed by an icosidodecahedron (ICD), an icosahedron (ICS), a dodecahedron (DO), and a tetrahedron (TE) at the center (Fig. 1a). In terms of crystallography, both 1/1 and 2/1 ACs are periodic crystals in 3D space, each with a distinct cubic space group, which leads to different Tsai-cluster arrangements in their unit cells (Fig. 1b) [15, 5]. The 1/1 AC typically forms a body-centered cubic (bcc) structure with the space group Im-3 (However, we noticed superstructure reflections for the Ce-Au-Al 1/1 AC which suggest a lower symmetry, see Supplementary Fig. S1). In contrast, the 2/1 AC adopts a primitive cubic structure with the space group Pa-3. The 16-atom polyhedron, constructed from the vertices of the ICS along the pseudo-5f axis, is considered the coordination environment for the Ce atoms (Fig. 1c,d).
@@
At high temperatures, above $\sim 10$ K the behavior is dominated by the phonons. At the lowest temperatures, below 0.3 K, the upturn in $C/T$ is due to a nuclear specific heat arising from 197Au with nuclear spin $I=3/2$ and 27Al with $I=5/2$ [21], apparent at high magnetic fields. The dominant contribution of the low-temperature specific heat resembles a magnetic Schottky anomaly, shifting up in temperature with increasing magnetic field.
@@
Figure 3: Specific heat and entropy of 1/1 AC and 2/1 AC. (a) and (c). The specific heat after removing the phonon contribution (see Supplementary Fig. S3 and S4) is shown as $C/T$ for 1/1 and 2/1 AC, respectively. (b) and (d). Corresponding entropy of the systems. The dotted gray line indicates the value $R$ln2. (e). Sketch of splitting of Ce3+ ($J$=5/2) degenerate levels, due to the non-cubic crystalline field, into three Kramers doublets. Crystal field disorder or magnetic interactions in the system could be the reason for lifting the Kramers spin degeneracy (LKSD). The magnetic field further splits the ground state doublet $|0\rangle$ as seen in the panel e.
Figure 4: Temperature and magnetic field dependence of the ac susceptibility. (a) and (b). The ac-susceptibility ($\chi$) for 1/1 and 2/1 AC for various magnetic fields in the log-lin scale. The inset shows the inverse susceptibility for several magnetic fields at low temperatures. The dotted line corresponds to $1/\chi\propto T$. (c). Calculated temperature dependence of the susceptibility for a two-level system for magnetic fields corresponding to panel a and b. (d) and (e). Magnetic field dependence of ac-susceptibility ($\chi$) for several temperatures in log-log scale. The applied magnetic field cuts off this diverging behavior and $\chi$ starts to decrease above the cut-off field ($B_{\alpha}$), shown as arrows. (f). Corresponding calculation for the field dependence of susceptibility for a two-level system for magnetic fields in panel d and e.
As we apply the magnetic field, the peak in $C/T$ shifts to higher temperature. This magnetic field behavior of the temperature dependence shows a maximum in $C$ that takes approximately the same value for all applied magnetic fields (see Supplementary Fig. S4). This indicates that the behavior is originating from a Schottky anomaly. The corresponding entropies, obtained as the integral of $C/T$, are shown in Fig. 3b and d. As seen from the entropy analysis, the entropy saturates to a value close to $\sim R\ln 2$ for both ACs at around 10 K. This indicates that the Schottky anomaly arises from a two-level system.
The non-cubic crystal field splits the Ce3+ ($J=5/2$) ground state into three Kramers doublets [22, 23, 24, 25]. A sketch of the energy level splitting is shown in Fig. 3e, where the ground state Kramers doublet is denoted as $|0\rangle$. The external magnetic field further splits the ground state doublet due to the Zeeman effect. This effect of the magnetic field on the Kramers doublet is seen in the specific heat in Fig. 3a and c. This ground state Kramers doublet, thus, determines the low-temperature behavior in specific heat.
@@
The presence of a significant zero-field specific heat indicates that the Kramers doublet ground state is not degenerate in the absence of magnetic field. This lifting of Kramers degeneracy could result from crystal field disorder [26] or magnetic interactions within the system [27, 28, 29, 30, 26].
@@
$W=\frac{\mu_{0}N_{\mathrm{A}}\mu^{2}}{3k_{\mathrm{B}}}$ (1)
@@
is the Curie constant, and $\mu=g\sqrt{J(J+1)}\mu_{\mathrm{B}}$ where $g$ is the Landé g-factor. Note that in this temperature range $\chi\sim 0.005$ emu/mol-Ce Oe (see Supplementary Fig. S5) which is significantly larger than any background from diamagnetic or electronic contributions [31, 32]. The Curie temperature extracted from a fit of $1/\chi$ vs $T$, shown in Supplementary Fig. S5, is $-10\pm 0.5\,\mathrm{K}$ and $-6\pm 1\,\mathrm{K}$ for 1/1 and 2/1 ACs, respectively, indicating weakly antiferromagnetic coupling. The effective magnetic moment $\mu_{\mathrm{exp}}$ of Ce obtained from this high-temperature region is $\mu_{\mathrm{exp}}=2.53\pm 0.02\,\mu_{\mathrm{B}}$ and $\mu_{\mathrm{exp}}=2.50\pm 0.02\,\mu_{\mathrm{B}}$ for 1/1 and 2/1 ACs, respectively (see Supplementary Fig. S5). This effective magnetic moment is close to the moment for free Ce3+ ion ($\mu_{\mathrm{cal}}=2.54\,\mu_{\mathrm{B}}$).
@@
Going to lower temperatures, the zero-field susceptibility no longer follows the high-temperature Curie-Weiss law. The insets of Fig. 4a and 4b show the low-temperature inverse susceptibility. A Curie-Weiss fit for temperatures between 2-10 K yields $\mu_{\mathrm{exp}}=1.98\pm 0.02\,\mu_{\mathrm{B}}$ and $\mu_{\mathrm{exp}}=2.10\pm 0.02\,\mu_{\mathrm{B}}$ for 1/1 and 2/1 ACs, respectively (see Supplementary Fig. S5), with Curie temperatures well below 1 K. The low-temperature zero-field $\chi$, thus, shows essentially a power-law behavior ${\chi\propto 1/T}$. The fairly large $\mu_{\mathrm{exp}}$ implies that the ground state Kramers doublet has a significant contribution of $m_{J}=5/2$.
@@
The diverging behavior in $\chi(T)$ is suppressed when an external magnetic field is applied, as seen in Fig. 4a and b. Figure 4d and 4e show the magnetic field dependence of 1/1 and 2/1 AC, respectively, for several temperatures. For a given temperature there exists a cutoff field $B_{\alpha}$ above which this suppression happens. Well below $B_{\alpha}$, the susceptibility becomes independent of the magnetic field and follows the zero-field $\chi(T)$. Above $B_{\alpha}$, the decrease of $\chi$ with field follows a power-law behavior. At the highest fields and lowest temperatures, the behavior becomes independent of temperature, while at low fields the behavior becomes independent of field. This indicates a temperature-field competition that governs the behavior of $\chi$.
The diverging zero-field susceptibility $\chi\propto 1/T$ could naively be attributed to a simple, non-interacting paramagnet with $\theta=0$. From the entropy in Fig. 3b and  3d we can see that the low temperature behavior of the system comes from the ground state Kramers doublet. The doublet acts as an effective spin 1/2 system, thus, $J_{\mathrm{eff}}=1/2$ at low temperatures [33, 34]. The ac-susceptibility of a $J_{\mathrm{eff}}=1/2$ Kramers doublet system in magnetic field is given by
$\chi_{\mathrm{sch}}=\frac{W_{\mathrm{K}}}{T}$ $\left[1-\tanh^{2}\left({\frac{g_{\mathrm{K}}\mu_{\mathrm{B}}B}{2k_{\mathrm{B}}T}}\right)\right]$ (2)
where
$W_{\mathrm{K}}=\frac{\mu_{0}Ng_{\mathrm{K}}^{2}\mu_{\mathrm{B}}^{2}}{4k_{\mathrm{B}}}.$ (3)
Note that the $g$-factor in Eq. (2) corresponds to the Kramers doublet and not the Landé $g$-factor. The zero field low-temperature fit of $1/\chi$ vs. $T$ (see Supplementary Fig. S5) directly gives $W_{\mathrm{K}}\approx 0.5$, which in turn gives $g_{\mathrm{K}}\approx 2.3$. The temperature and field dependence of the susceptibility for the ground state Kramers doublet from Eq. (2) using $g_{\mathrm{K}}=2.3$ and $W_{\mathrm{K}}=0.5$ (emu K/Oe)/mol is shown in Fig. 4c and 4d for the corresponding fields and temperatures.
@@
For a quantum critical system, the susceptibility has a form of ${1/\chi\propto T^{\beta}}$, where $\beta$ is a critical exponent. This critical exponent characterizes the spectrum of critical fluctuations in a quantum critical system [35]. For Ce-Au-Al ACs at low temperatures (2-10 K), this critical exponent seems to be approximately equal to one ($\beta=0.94\pm 0.01$ for 1/1 AC and $0.97\pm 0.01$ for 2/1 AC). This suggests that $\beta\approx 1$ and most of the behavior seen in ac-susceptibility comes from the Zeeman splitting of the Kramers doublet in magnetic fields similar to the behavior seen in specific heat.
From transport measurements [13], it is seen that the ACs of the Ce-Au-Al show a Kondo resonance peak at $T_{\mathrm{K}}\approx$1 K. This indicates that the interaction between the conduction electrons and $f$-electrons becomes important only below 1 K. The conduction electron effective mass above 5 K as extracted from the Sommerfeld coefficient evaluated in Supplementary Fig. S3 is rather modest, with $\gamma\approx 4$ mJ/mol-at.K2. Thus, the system does not display any significant heavy fermion characteristics, consistent with $T_{\mathrm{K}}$ being small.
The temperature-field competition is one of the defining features of quantum criticality in a broad range of systems [36, 35, 37, 38, 39, 40] including the quantum critical quasicrystal Yb-Au-Al [11, 12, 41]. However, the susceptibility of even a non-interacting two-level system behaves very similarly in magnetic fields, as seen from Figs. 4c and 4f. There are differences, though. The measured $\chi$ approximately follows a power-law behavior in high fields, which is an indication of quantum critical behavior, whereas the $\chi_{\mathrm{sch}}$ calculations appear as a more sudden drop above the cutoff field.
Most of the heavy fermion systems have a Kramers doublet ground state [42, 43]. The strong interaction of the conduction electrons with the $f$-electrons induces the quantum critical behavior observed in these systems [42, 43]. In heavy fermion quantum critical systems, the critical fluctuations modify this behavior of the Kramers doublet by changing the $\beta$ and significantly reducing the $g_{\mathrm{K}}$ of the doublet. Indications of interactions in the Ce-Au-Al approximants studied here are seen both at zero field at at high field. As seen from $C/T$, the Kramers degeneracy is lifted even in zero magnetic field. At high fields, the power-law dependence of $\chi$ gives a weaker field dependence than expected from a non-interacting system.
From the present data, it is not clear whether the ground state of the system is magnetically ordered or a paramagnetic state. If magnetic ordering sets in, the transition temperature is below about 0.2 K, since no clear signatures are seen in specific heat. The reason for this could be the geometric frustration from the ACs structure. The complex atomic clusters allow multiple rare-earth sites with different crystal fields and orientation with respect to the applied field. This likely contributes to the suppression of long-range magnetic order in the system.